Determine the numerical apperature of a step index fibre when the core refractive index n1=1.5 and cladding refractive index n2=1.48 find the maximum angle for entrance of light if the fibre is placed in the air
1
Expert's answer
2020-04-05T14:43:38-0400

Numerical aperture (NA) ="\sqrt{n_1^2-n_2^2}=\sqrt{(1.5)^2-(1.48)^2}=\sqrt{2.25-2.19}=0.245"